OpenStudy (e.cociuba):

Solve the equation using the quadratic formula: x^2 + 4x +3 = 0

OpenStudy (anonymous):

umm.. maybe (x+3)(x+1)=0 so, x=-3 or x=-1

OpenStudy (e.cociuba):

omg tht so makes sense! Thanks! Can u help w/ my las one too? 4b^2 + 8b +7= 4

OpenStudy (e.cociuba):

Need help w/ this one: 4b^2 + 8b +7= 4

OpenStudy (mathstudent55):

4b^2 + 8b +7= 4 4b^2 + 8b + 3 = 0 4x^2 + 8x + 3 = 0